Jim’s salary is 85% of Karen’s. Together they made $222 last Saturday. How much did Jim earn that day.
Thank you!

Answer by Theo(13342)

let j = jim's salary and let k = karen's salary.
j + k = 222
j = .85 * k
replace j with .85 * k to get:
.85 * k + k = 222
combine like terms to get:
1.85 * k = 222
solve for k to get:
k = 222/1.85 = 120
j = 222 - 120 = 102
102/120 = .85

jim's salary is 102 that day.
karen's salary is 120 that day.
jim's salary is 102 / 120 = .85 * 100 = 85% of karen's that day.